OpenStudy (anonymous):

Solve the following for all real values of x. 2/x+1=x-2/2 5/e^x +1=1

OpenStudy (nurali):

\[\frac{ 2 }{ x+1 }=\frac{ x- 2}{ 2 }\] or \[\frac{ 2 }{ x }+1=\frac{ x-2 }{ 2 }\]

OpenStudy (nurali):

Which one?

OpenStudy (nurali):

reply @peanut1995

OpenStudy (anonymous):

first one

OpenStudy (nurali):

ok

OpenStudy (nurali):

\[\frac{ 2 }{ x+1 }=\frac{ x-2 }{ 2 }\] Cross Multiply \[4=(x+1)(x-2)\] \[4=x^2-2x+x-2\] \[4=x^2-x-2\] \[x^2-x-6=0\] Factorize \[(x+2)(x-3)=0\] \[x=-2,3\]

OpenStudy (nurali):

\[\frac{ 5 }{ e^x+1 }=1\] or \[\frac{ 5 }{ e^x }+1=1\]

OpenStudy (nurali):

Which one?

OpenStudy (anonymous):

first one

OpenStudy (nurali):

ok

OpenStudy (nurali):

\[\frac{ 5 }{ e^x+1 }=1\] Cross Multiply \[5=e^x+1\] \[e^x=5-1\] \[e^x=4\] Take the natural log of both sides: Note: ln[e] = 1 e^x = 4 ln[e^x] = ln[4] x * ln[e] = ln[4] x = ln[4] x ≈ 1.386

OpenStudy (anonymous):

okay would you mind helping me on one more?

OpenStudy (nurali):

i will try.

OpenStudy (anonymous):

\[\sqrt{x-1} -5/\sqrt{x-1}=0\]

OpenStudy (nurali):

\[\sqrt{x-1}-\frac{ 5 }{\sqrt{x-1} }=0\] Taking LCM \[\frac{ \sqrt{x-1}\sqrt{x-1}-5 }{ \sqrt{x-1} }=0\] Cross Multiply \[x-1-5=0\] \[x-6=0\] \[x=6\]

OpenStudy (anonymous):

Thank you so so much:) I really appreciate it

OpenStudy (nurali):

My Pleasure.